# Solving cluster moment relaxation with hierarchical matrix

### v1

@@ 2 Proposed convex relaxations
We now use these four ingredients to provide a convex relaxation for the energy minimization problem in terms of the moment matrix $M\in\mathbb{C}^{(CK+1)\times(CK+1)}$
@@
Here, the expectations are taken entry-wise and defined as $\mathbb{E}_{\mu}(\cdot):=\int(\cdot)\mu(dx)$ and $\mathbb{E}_{\rho}(\cdot):=\langle\cdot,\rho\rangle$. In what follows, for a matrix $A\in\mathbb{C}^{(CK+1)\times(CK+1)}$, we partition $A$ into
$A:=\begin{bmatrix}A^{(2)}&A^{(1)}\\ {A^{(1)}}^{*}&A^{(0)}\end{bmatrix},\quad A^{(2)}\in\mathbb{C}^{CK\times CK},\ A^{(1)}\in\mathbb{C}^{CK\times 1},\ M^{(0)}\in\mathbb{C}.$ (6)
@@
$A^{(2)}=[A^{(2)}_{ij}]_{i=1,j=1}^{K}$ (7)
where each $A^{(2)}_{ij}$ is a $C\times C$ block.
@@ 2.1.1 Cluster moment relaxation for the TFI
$\mathbf{v}:=[\mathbf{v}_{i}]_{i=1}^{N},\quad\mathbf{v}_{i}=\begin{bmatrix}{\sigma_{i}^{x}}^{*}&{\sigma_{i}^{y}}^{*}&{\sigma_{i}^{z}}^{*}\end{bmatrix}^{*},$ (17)
@@
$M^{(2)}(i,j)=\mathrm{Tr}(\mathbf{v}(i){\mathbf{v}(j)}^{*}\rho),\quad M^{(1)}(i)=\mathrm{Tr}(\mathbf{v}(i)\rho).$ (18)
@@ 3 Standard augmented Lagrangian method
($\mathcal{F}^{*}(\{\Lambda_{ij}\}_{i<j},\{\lambda_{j}\}_{j},\gamma):=\begin{bmatrix}\mathcal{D}^{*}(\lambda_{1})&\mathcal{A}^{*}_{U}(\Lambda_{12})&\cdots&\mathcal{A}^{*}_{U}(\Lambda_{1K})&&\mathcal{D}_{U}^{*}(\lambda_{1})\\ \mathcal{A}^{*}_{L}(\Lambda_{12})&\mathcal{D}^{*}(\lambda_{2})&\ddots&&&\vdots\\ \vdots&\ddots&\ddots&\mathcal{A}^{*}_{U}(\Lambda_{(K-1)K})&&\\ \mathcal{A}^{*}_{L}(\Lambda_{1K})&\cdots&\mathcal{A}^{*}_{L}(\Lambda_{(K-1)K})&\mathcal{D}^{*}(\lambda_{K})&&\mathcal{D}_{U}^{*}(\lambda_{K})\\ \mathcal{D}_{L}^{*}(\lambda_{1})&\cdots&&\mathcal{D}_{L}^{*}(\lambda_{K})&&\gamma\end{bmatrix}.$ (26)
@@
In Algorithm 1, it is worth noting that the primal variable $M$ always satisfies the linear constraint in (P) during the updates. This arises from the property that ALM always provides dual-feasible variables [8] (where (P) is the dual problem of (D)).
@@ 4.1.1 Validity of the hierarchical matrix representation for $S$
$\min_{y,z}\ \ \text{Err}_{S}^{2}:=\frac{\|\mathcal{H}(y,t)-S^{*}\|_{F}^{2}}{\|S^{*}\|_{F}^{2}},$ (34)
by running the limited-memory BFGS algorithm provided in the Manopt toolbox [6] for $1000$ iterations. The approximation errors are presented in Table 2. From the table, we can see that even with fixed $r_{l}=20$ for all $1\leq l\leq m$, we obtained similar accuracy for different system sizes. Therefore, we assume one can use a fixed rank approximation in the hierarchical matrix even for large system sizes.
@@ 5 ALM with hierarchical primal and dual PSD variables
$\begin{split}&\mathrm{argmin}_{\tilde{M}}\quad\|\tilde{M}-\big{(}M+\sigma\left(S-J+\mathcal{F}^{*}(\{\Lambda_{ij}\}_{i<j},\{\lambda_{j}\}_{j},\gamma)\right))\big{)}\|_{F}^{2},\\ &\hskip 14.22636pt\text{s.t.}\hskip 42.67912pt\mathcal{A}(\tilde{M})=b.\end{split}$ (41)
Here, $S$ is represented hierarchically as $S=\mathcal{H}(y,t)$, with $\{\Lambda_{ij}\}_{i<j},\{\lambda_{j}\}_{j},\gamma$ being defined in (28), (29) and (30). The linear constraint $\mathcal{A}(\tilde{M})=b$ summarizes all linear constraints on the matrix $M$. Although this constraint in (41) may seem redundant since the primal variable always satisfies the linear constraint in (P) during the updates, it becomes crucial when replacing $\tilde{M}$ with a compressed representation, as the constraint will not be naturally satisfied.
To this end, we aim to find an affine function $g:\mathbb{H}^{CK+1}\rightarrow\mathbb{H}^{CK+1}$ such that $\mathcal{A}(g(A))=b$ for any $A\in\mathbb{H}^{CK+1}$. A choice of $g$ is such that $g(A)=\mathcal{P}_{\text{null}(\mathcal{A})}(A)+{\mathcal{A}}^{\dagger}(b)$. Here, ${\mathcal{A}}^{\dagger}$ denotes the pseudo-inverse of the operator $\mathcal{A}.$ Moreover, the linear constraints in cluster moment relaxation are local in nature and only constrain local blocks. This means one can choose $g$ with
$g(A):=\begin{bmatrix}g^{(1)}(A^{(2)}_{11},A^{(1)}_{1})&&&&&\tilde{g}^{(1)}(A^{(2)}_{11},A^{(1)}_{1})\\ &g^{(1)}(A^{(2)}_{22},A^{(1)}_{2})&&&&\vdots\\ &&\ddots&&&\\ &&&g^{(1)}(A^{(2)}_{KK},A^{(1)}_{K})&&\tilde{g}^{(1)}(A^{(2)}_{KK},A^{(1)}_{K})\\ \tilde{g}^{(1)}(A^{(2)}_{11},A^{(1)}_{1})^{*}&\cdots&&\tilde{g}^{(1)}(A^{(2)}_{KK},A^{(1)}_{K})^{*}&&g^{(0)}(A^{(0)})\end{bmatrix}+\\ \begin{bmatrix}0_{C\times C}&g^{(2)}(A^{(2)}_{12},A^{(2)}_{21})&\cdots&g^{(2)}(A^{(2)}_{1K},A^{(2)}_{K1})&&0_{C\times 1}\\ g^{(2)}(A^{(2)}_{12},A^{(2)}_{21})^{*}&0_{C\times C}&\ddots&&&\vdots\\ \vdots&\ddots&\ddots&g^{(2)}(A^{(2)}_{(K-1)K},A^{(2)}_{K(K-1)})&&\\ g^{(2)}(A^{(2)}_{1K},A^{(2)}_{K1})^{*}&\cdots&g^{(2)}(A^{(2)}_{(K-1)K},A^{(2)}_{K(K-1)})^{*}&0_{C\times C}&&0_{C\times 1}\\ 0_{1\times C}&\cdots&&0_{1\times C}&&0\end{bmatrix}.$ (42)
where $g^{(2)}:{\mathbb{C}^{C\times C}}\times\mathbb{C}^{C\times C}\rightarrow\mathbb{C}^{C\times C}$, $g^{(1)}:{\mathbb{C}^{C\times C}}\times\mathbb{C}^{C\times 1}\rightarrow\mathbb{C}^{C\times C}$, $\tilde{g}^{(1)}:{\mathbb{C}^{C\times C}}\times\mathbb{C}^{C\times 1}\rightarrow\mathbb{C}^{C\times 1}$ and $g^{(0)}:\mathbb{C}\rightarrow\mathbb{C}$ are all affine transformations that depend on the constant vector $b$. For example, in order to satisfy the linear constraints of the TFI model of $N$ sites (Section 2.1.1), we let $g:\mathbb{H}^{3N+1}\rightarrow\mathbb{H}^{3N+1}$ be:
$g(A)=I_{3N+1}+\real\Bigg{(}\begin{pmatrix}0&A^{(2)}_{12}&\cdots&A^{(2)}_{1N}&A^{(1)}_{1}\\ A^{(2)}_{21}&0&\cdots&A^{(2)}_{2N}&A^{(1)}_{2}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ A^{(2)}_{N1}&A^{(2)}_{N2}&\cdots&0&A^{(1)}_{N}\\ {A^{(1)}_{1}}^{*}&{A^{(1)}_{2}}^{*}&\cdots&{A^{(1)}_{N}}^{*}&0\end{pmatrix}\Bigg{)}\\ +2\begin{pmatrix}\mathcal{B}^{*}({i\mkern 1.0mu}\real(A^{(1)}_{1}))&&&\\ &\ddots&&\\ &&\mathcal{B}^{*}({i\mkern 1.0mu}\real(A^{(1)}_{N}))&\\ &&&0\end{pmatrix},$ (43)
with
$\mathcal{B}(D):=\begin{pmatrix}D(2,3)\\ D(3,1)\\ D(1,2)\end{pmatrix},\quad\mathcal{B}:\mathbb{H}^{3}\rightarrow\mathbb{C}^{3}.$
This reparameterization of primal-feasible matrices allows us to rewrite (41) in an equivalent form:
$\begin{split}&\mathrm{argmin}_{A\in\mathbb{H}^{CK+1}}\quad\|g(A)-\big{(}M+\sigma\left(S-J+\mathcal{F}^{*}(\{\Lambda_{ij}\}_{i<j},\{\lambda_{j}\}_{j},\gamma)\right)\big{)}\|_{F}^{2}.\end{split}$ (44)
Our final step is to approximate $A\in\mathbb{H}^{CK+1}$ with a hierarchical matrix to reduce the cost of the loss function evaluation:
$\begin{split}&\mathrm{argmin}_{x,v,A}\quad\|g(A)-\big{(}M+\sigma\left(S-J+\mathcal{F}^{*}(\{\Lambda_{ij}\}_{i<j},\{\lambda_{j}\}_{j},\gamma)\right)\big{)}\|_{F}^{2}\\ &\hskip 14.22636pt\text{s.t.}\hskip 99.58464ptA=\mathcal{H}(x,v),\end{split}$ (45)
where $x=\{x^{(i)}\in\mathbb{C}^{CK\times r_{i}}\}_{1\leq i\leq m}$ and $v\in\mathbb{C}^{(CK+1)\times 1}$ for some pre-specified number of levels $m$, and number of columns $r_{1},\cdots,r_{m}$. In other words, we propose a hierarchical representation for the parameterization of any primal-feasible variable $M$. This allows efficient evaluation of the loss function in (45). If the current guess $M$ is represented as $M=g(\mathcal{H}(x,v))$ for some $x$ and $v$, and $S=\mathcal{H}(y,t)$, then the main computational task in evaluating the loss function (45) can be reduced.
@@
1:$y,t$ for the dual variable $S=\mathcal{H}(y,t)$, $x,v$ for the primal variable $M=g(\mathcal{H}(x,v))$, and penalty parameter $\sigma>0$
@@
3:     $M\hskip 2.84526pt\leftarrow g(\mathcal{H}(x,v))$ where $g$ is defined in (42)
@@ 5.1 Validity of the hierarchical matrix representation for M
In this section, we investigate the validity of representing $M$ by a hierarchical matrix $\mathcal{H}(x,v)$, using the same method as in section 4.1.1 for the dual PSD variable $S$. Instead of fitting the dual PSD variable $S$, we fitted the primal PSD variable $M$ resulting from solving (D) with the Hamiltonian specified in (16), for system sizes $N=64,128,256$, an external magnetic field strength parameter $h=1$ and an accuracy of $10^{-4}$. Let $M^{*}$ be the approximate solution. Let $x,v$ be complex-valued parameters that parameterize a hierarchical matrix of the corresponding size, with the number of levels $m=3,4,5$ respectively for $N=64,128,256$. The number of columns for each level was fixed at $20$ for all system sizes. We solved the following optimization problem with system sizes $N=64,128,256$ for the variables $x$ and $t$:
@@
by running the limited-memory BFGS algorithm provided in the Manopt toolbox [6] for $1000$ iterations. The approximation errors are presented in Table 3. The experiment results indicate that as the system size increases, the relative error of the fitted primal PSD variable also tends to increase. This likely contributes to the failure of Algorithm 4 for large system sizes, such as $N=4096$ (see Section 6 for further details on the experiments).
@@ 6 Numerical experiments
In this section, we present numerical experiments for the $1$-D TFI model using Algorithms 3 and 4, with system sizes $N\in\{64,128,256,512,1024,2048,4096\}$. The penalty parameter is initialized at $\sigma=0.1$ and is adjusted dynamically [25] based on primal and dual feasibility to speed up the convergence of the ALM algorithm. For both algorithms, the number of levels $m$ in the hierarchy is set to be $m=3,4,\cdots,9$ for $N=64,128,\cdots,4096$, and the number of columns for all levels is set to be $r=20$. In Algorithm 3, $y$, $t$ and $M$ are initialized from the standard normal distribution, with $M$ being primal-feasible. This is achieved by having the initial primal variable $M=g(A)$ for some $A\in\mathbb{H}^{3N+1}$, where $g$ is defined in (43). In Algorithm 4, $x$, $v$, $y$ and $t$ are randomly initialized from the standard normal distribution. Throughout the updates of Algorithms 3 and 4, we evaluate the accuracy of approximate solutions by monitoring the relative primal feasibility, the relative dual feasibility, and the relative duality gap, as detailed in the following.
The primal feasibility of the variable $M$ is governed by how well it satisfies the PSD and linear constraints in (P). In Algorithm 3, $M$ is directly updated as a dense matrix, while in Algorithm 4, $M$ is maintained using a compressed representation that satisfies the linear constraints. Since $M$ is guaranteed to satisfy the linear constraints in (P), we only monitor its PSDness using the following measure:
$\eta_{P}:=\frac{\max(0,-\lambda_{\text{min}})}{1+\max(0,\lambda_{\text{max}})},$
where $\lambda_{\text{min}}$ and $\lambda_{\text{max}}$ are the smallest and largest eigenvalues of $M$.
@@
We terminate the algorithm when $\eta:=\max(\eta_{P},\eta_{D},\eta_{g})\leq 10^{-3}$, or when the ALM algorithm has run for $1500$ iterations. It is important to highlight that we exploit the hierarchical structure present in the PSD primal and dual variables to efficiently evaluate these convergence metrics.
@@
The relative errors for Algorithms 3 and 4 are given in Tables 4 and 5. Additionally, we present the evolution of our convergence metrics as a function of the ALM iteration number in Figures 2 and 3, focusing on the $1$-D TFI model with a fixed external magnetic field strength parameter $h=1$ and various system sizes. Alongside the relative primal and dual feasibility measures and the relative duality gap, we also track the per-site primal objective change between subsequent iterations. All convergence metrics are transformed using a base-$10$ logarithm function. Within $1500$ ALM iterations, all metrics drop below $10^{-3}$ for experiments with an external magnetic field strength parameter $h=1$. We note that it is harder to make Algorithm 4 converge compared to Algorithm 3. For example, we are not able to get the case for $N=4096$ to converge to the prescribed accuracy $10^{-3}$. We suspect this is due to the fact that it is harder to fit the primal PSD variable by a low-rank hierarchical matrix compared to the dual PSD variable.

### v2

@@ 2 Proposed convex relaxations
The meaning of item 3 and 4 will be illustrated through an example in the next subsection. We now use these four ingredients to provide a convex relaxation for the energy minimization problem in terms of the moment matrix $M\in\mathbb{C}^{(CK+1)\times(CK+1)}$
@@
Here, the expectations are taken entry-wise and defined as $\mathbb{E}_{\mu}(\cdot):=\int(\cdot)\mu(dx)$ and $\mathbb{E}_{\rho}(\cdot):=\langle\cdot,\rho\rangle$. In what follows, for any matrix $A\in\mathbb{C}^{(CK+1)\times(CK+1)}$, we partition $A$ into
$A:=\begin{bmatrix}A^{(2)}&A^{(1)}\\ {A^{(1)}}^{*}&A^{(0)}\end{bmatrix},\quad A^{(2)}\in\mathbb{C}^{CK\times CK},\ A^{(1)}\in\mathbb{C}^{CK\times 1},\ A^{(0)}\in\mathbb{C}.$ (6)
@@
$A^{(2)}=[A^{(2)}_{ij}]_{i=1,j=1}^{K},\quad A^{(1)}=[A_{i}^{(1)}]_{i=1}^{K}$ (7)
where each $A^{(2)}_{ij}$ is a $C\times C$ block, and each $A_{i}^{(1)}$ is a $C\times 1$ vector. For example, $M^{(2)}=\mathbb{E}_{\mu}(\mathbf{v}\mathbf{v}^{*})$ or $M^{(2)}=\mathbb{E}_{\rho}(\mathbf{v}\mathbf{v}^{*})$.
@@
$J$ is a cost matrix that encodes the information of $\mathcal{\epsilon}$ and $\hat{H}$ in (1) and (2) respectively. The precise meaning of the terms in (8) such as $J,w,z$ are illustrated in the next subsection through the example of a quantum spin model.
@@ 2.1.1 Cluster moment relaxation for the TFI
${\color[rgb]{0,0,1}\mathbf{v}:=[\mathbf{v}_{i}]_{i=1}^{N},\quad\mathbf{v}_{i}=\begin{bmatrix}{\sigma_{i}^{x}}^{*}&{\sigma_{i}^{y}}^{*}&{\sigma_{i}^{z}}^{*}\end{bmatrix}^{*},}$ (17)
@@
$M^{(2)}_{ij}=\mathrm{Tr}(\mathbf{v}_{i}{\mathbf{v}_{j}}^{*}\rho),\quad M^{(1)}_{i}=\mathrm{Tr}(\mathbf{v}_{i}\rho).$ (18)
@@
For the precise definitions of $\mathcal{A}_{U}$, $\mathcal{A}_{L}$, $\mathcal{D}_{U}$, $\mathcal{D}_{L}$, $\mathcal{D}$, and $J,w,z$, we refer readers to the appendix.
@@ 3 Standard augmented Lagrangian method
($\mathcal{F}^{*}(\{\Lambda_{ij}\}_{i<j},\{\lambda_{j}\}_{j},\gamma):=\begin{bmatrix}\mathcal{D}^{*}(\lambda_{1})&\mathcal{A}^{*}_{U}(\Lambda_{12})&\cdots&\mathcal{A}^{*}_{U}(\Lambda_{1K})&&\mathcal{D}_{U}^{*}(\lambda_{1})\\ \mathcal{A}^{*}_{L}(\Lambda_{12})&\mathcal{D}^{*}(\lambda_{2})&\ddots&&&\vdots\\ \vdots&\ddots&\ddots&\mathcal{A}^{*}_{U}(\Lambda_{(K-1)K})&&\\ \mathcal{A}^{*}_{L}(\Lambda_{1K})&\cdots&\mathcal{A}^{*}_{L}(\Lambda_{(K-1)K})&\mathcal{D}^{*}(\lambda_{K})&&\mathcal{D}_{U}^{*}(\lambda_{K})\\ \mathcal{D}_{L}^{*}(\lambda_{1})&\cdots&&\mathcal{D}_{L}^{*}(\lambda_{K})&&\gamma\end{bmatrix}$ (26)
is hermitian so we have $\mathcal{A}_{U}^{*}(\Lambda_{ij})=\left(\mathcal{A}_{L}^{*}(\Lambda_{ij})\right)^{*}$ for $i<j$, and $\mathcal{D}^{*}(\lambda_{j})=\left(\mathcal{D}^{*}(\lambda_{j})\right)^{*},\,\mathcal{D}_{U}^{*}(\lambda_{j})=\left(\mathcal{D}_{L}^{*}(\lambda_{j})\right)^{*}$ for $j\in[K]$.
@@
In Algorithm 1, it is worth noting that the primal variable $M$ always satisfies the linear constraint in (P) during the updates. This arises from the property that ALM always provides dual-feasible variables [alm_dual_feasible] (where (P) is the dual problem of (D)).
@@ 4.1.1 Validity of the hierarchical matrix representation for $S$
$\min_{y,t}\ \ \text{Err}_{S}^{2}:=\frac{\|\mathcal{H}(y,t)-S^{*}\|_{F}^{2}}{\|S^{*}\|_{F}^{2}},$ (34)
by running the limited-memory BFGS algorithm provided in the Manopt toolbox [6] for $100$ iterations. The approximation errors are presented in Table 2. From the table, we can see that even with fixed $r_{l}=20$ for all $1\leq l\leq m$, we obtained similar accuracy for different system sizes. Therefore, we assume one can use a fixed rank approximation in the hierarchical matrix even for large system sizes.
@@ 5 ALM with hierarchical primal and dual PSD variables
$\begin{split}&\mathrm{argmin}_{\tilde{M}}\quad\|\tilde{M}-\big{(}M+\sigma\left(S-J+\mathcal{F}^{*}(\{\Lambda_{ij}\}_{i<j},\{\lambda_{j}\}_{j},\gamma)\right))\big{)}\|_{F}^{2},\\ \end{split}$ (41)
Here, $S$ is represented hierarchically as $S=\mathcal{H}(y,t)$, with $\{\Lambda_{ij}\}_{i<j},\{\lambda_{j}\}_{j},\gamma$ being defined in (28), (29) and (30). For simplicity, we still use (25) to represent the primal variable $M$ as a hierarchical matrix, i.e. $M=\mathcal{H}(x,v)$, where $x=\big{\{}x^{(1)},\cdots,x^{(m)}\big{\}}$ is a set of matrices with $x^{(l)}\in\mathbb{C}^{CK\times r_{l}}$, and $v\in\mathbb{C}^{CK+1}$. Then (41) becomes:
$\begin{split}&\mathrm{argmin}_{x,v}\quad\|\mathcal{H}(x,v)-\big{(}M+\sigma\left(S-J+\mathcal{F}^{*}(\{\Lambda_{ij}\}_{i<j},\{\lambda_{j}\}_{j},\gamma)\right))\big{)}\|_{F}^{2},\\ \end{split}$ (42)
@@
1:$y,t$ for the dual variable $S=\mathcal{H}(y,t)$, $x,v$ for the primal variable $M=\mathcal{H}(x,v)$, and penalty parameter $\sigma>0$
@@
3:     $M\hskip 2.84526pt\leftarrow\mathcal{H}(x,v)$
@@ 5.1 Validity of the hierarchical matrix representation for M
In this section, we investigate the validity of representing $M$ by a hierarchical matrix $\mathcal{H}(x,v)$, using the same method as in section 4.1.1 for the dual PSD variable $S$. Instead of fitting the dual PSD variable $S$, we fitted the primal PSD variable $M$ resulting from solving (D) with the Hamiltonian specified in (16), for system sizes $N=64,128,256$, an external magnetic field strength parameter $h=1$ and an accuracy of $10^{-4}$. Let $M^{*}$ be the approximate solution. Let $x,v$ be complex-valued parameters that parameterize a hierarchical matrix of the corresponding size, with the number of levels $m=3,4,5$ respectively for $N=64,128,256$. The number of columns for each level was fixed at $20$ for all system sizes. We solved the following optimization problem with system sizes $N=64,128,256$ for the variables $x$ and $v$:
@@
by running the limited-memory BFGS algorithm provided in the Manopt toolbox [6] for $100$ iterations. The approximation errors are presented in Table 3. The experiment results indicate that as the system size increases, the relative error of the fitted primal PSD variable slowly increase, showing that the hierarchical matrix representation for $M$ can be valid for larger system sizes.
@@ 6 Numerical experiments
In this section, we present numerical experiments for the $1$-D TFI model using Algorithms 3 and 4, with system sizes $N\in\{64,128,256,512,1024,2048,4096\}$. The penalty parameter is initialized at $\sigma=0.1$ for $N\geq 1024$, $\sigma=1$ for $N\leq 512$ and is adjusted dynamically [24] based on primal and dual feasibility to speed up the convergence of the ALM algorithm. For both algorithms, the number of levels $m$ in the hierarchy is set to be $m=3,4,\cdots,9$ for $N=64,128,\cdots,4096$, and the number of columns for all levels is set to be $r=20$. In Algorithm 3, $y$, $t$ are initialized from the standard normal distribution, and $M_{0}=I_{3N+1}$. In Algorithm 4, $x$, $v$, $y$ and $t$ are randomly initialized from the standard normal distribution. Throughout the updates of Algorithms 3 and 4, we evaluate the accuracy of approximate solutions by monitoring the relative primal feasibility, the relative dual feasibility, and the relative duality gap, as detailed in the following.
The primal feasibility of the variable $M$ is governed by how well it satisfies the PSD and linear constraints in (P). In Algorithm 3, $M$ is directly updated as a dense matrix, while in Algorithm 4, $M$ is maintained using a compressed representation that satisfies the linear constraints. Since in Algorithm 3, $M$ is guaranteed to satisfy the linear constraints in (P), we only monitor its PSDness using the following measure:
$\eta_{P_{3}}:=\frac{\max(0,-\lambda_{\text{min}})}{1+\max(0,\lambda_{\text{max}})},$
where $\lambda_{\text{min}}$ and $\lambda_{\text{max}}$ are the smallest and largest eigenvalues of $M$. In Algorithm 4, we update $M$ by solving (34), which guaranteeds that $M$ is PSD. Therefore, we only monitor the linear constraint satisfaction by:
$\eta_{P_{4}}:=\frac{\|\mathcal{A}(M)-b\|_{2}}{1+\|b\|_{2}},$
[…]
@@
We terminate the algorithm when $\eta:=\max(\eta_{P},\eta_{D},\eta_{g})\leq 10^{-3}$, or when the ALM algorithm has run for $150$ iterations. It is important to highlight that we exploit the hierarchical structure present in the PSD primal and dual variables to efficiently evaluate these convergence metrics.
@@
The relative errors for Algorithms 3 and 4 are given in Tables 4 and 5. Additionally, we present the evolution of our convergence metrics as a function of the ALM iteration number in Figures 2 and 3, focusing on the $1$-D TFI model with a fixed external magnetic field strength parameter $h=1$ and various system sizes. Alongside the relative primal and dual feasibility measures and the relative duality gap, we also track the per-site primal objective change between subsequent iterations. All convergence metrics are transformed using a base-$10$ logarithm function. Within $150$ ALM iterations, all metrics drop below $10^{-3}$ for experiments with an external magnetic field strength parameter $h=1$.